# A Clinically-Grounded Two-Stage Framework for Renal CT Report Generation

### v1

@@ 4.2.3 CT Slice
From the 3D CT volumes, we extracted 2D slices explicitly referenced in the radiology reports (e.g., “image 45 series 4”). Slice positions were determined by sorting the DICOM series using the SliceLocation attribute, followed by index-based matching to the reported image number. Each slice was then aligned with its corresponding sentence and renal feature labels to form complete data triplets (report sentence, label, CT slice). For detection models, we retained the original Hounsfield Unit (HU) values and deferred augmentation to the training stage, treating augmentation parameters as part of the hyperparameter search. For generative models, we applied fixed windowing (window width = 400, window level = 50) and converted the slices to PNG format. All CT slices were manually reviewed to ensure they corresponded to the renal region. Slices that were misreferenced or misaligned by the radiologist were excluded to ensure high-quality inputs for downstream modeling.
@@ 4.3 Data Splitting and Label Balance
We adopted a 5-fold cross-validation strategy to ensure robustness and generalizability. Data splitting was performed at the annotation level, where each sample corresponds to a unique triplet of (report sentence, label, CT slice). This approach enables stratification across a larger number of samples while preserving label diversity within each fold. To mitigate sampling bias, we applied stratified sampling [28, 29] based on renal feature distributions. For highly imbalanced features—such as the “Exophytic” category, which included only one “endophytic” case—we manually ensured that the minority class was represented in both training and validation sets. The detailed label distribution for one representative fold is provided in Appendix B.
@@ 4.4 Model
To align with the radiology workflow and enhance clinical interpretability, we adopt a two-stage modeling pipeline (Figure 3). In the first stage, we extract renal abnormality features from 2D CT slices using tailored computer vision models. In the second stage, these features—along with the corresponding slice—are fed into a fine-tuned vision-language model (VLM) to generate a free-text report sentence. This modular structure enables both structured feature evaluation and flexible language generation.
4.4.1 Feature Extraction Model
We use a ResNet-34 backbone [30] as the base encoder for CT slice-level feature extraction. Each renal feature label is treated as a separate task:
@@ 4.4.1 Feature Extraction Model
The Exophytic and Tumor categories are extremely imbalanced in our dataset (e.g., only one endophytic case among 130 lesions), rendering supervised classification ineffective. For these categories, we employ anomaly detection techniques [31] trained exclusively on negative examples. All models are trained and evaluated using stratified 5-fold cross-validation (Section 4.3), with task-appropriate metrics such as AUROC, MAE, and F1 score.
@@ 4.4.2 Report Generation Model
We fine-tune Qwen-VL 2.5 [19], a vision-language model capable of interpreting paired image–text inputs. The generation prompt consists of two components: (1) a structured feature template populated with predicted renal feature labels, and (2) the corresponding CT slice. Together, these inputs are used to generate a radiology report sentence aligned with the annotated source. Model fine-tuning is performed using LLAMA-Factory [32] with instruction-style supervision. For each training sample, the input includes the structured prompt and CT slice, and the target is the expert-written report sentence. Prompt templates used for report generation are provided in Appendix A.
Figure 3: Overview of our two-stage pipeline for renal CT report generation. The Feature Extraction Model identifies structured renal feature labels from 2D CT slices, which are then used by the Report Generation Model to produce the final radiology report sentence.
@@ 4.5 Data Augmentation
To improve model robustness and leverage spatial continuity in CT volumes, we applied augmentation strategies inspired by prior work [12]. Specifically, for each annotated 2D CT slice, we included its adjacent slices at offsets of 11-1- 1 and +11+1+ 1, effectively tripling the training data and introducing slight anatomical variation without altering the underlying pathology. This augmentation was applied exclusively to the feature extraction model, and not to the report generation model, as the latter operates on individual CT slices with precise spatial references. We also standardized the spatial resolution of all CT slices by resizing each to a fixed dimension of $512\times 512512 × 512$ pixels. For slices larger than this size, we applied center cropping; for smaller ones, we padded the surrounding region with zeros. This resizing strategy ensured consistency across inputs while preserving lesion scale and spatial integrity—particularly important for size-sensitive tasks.
In addition to spatial normalization, we applied standard intensity preprocessing. Raw CT scans were stored in Hounsfield Units (HU), spanning a wide dynamic range that includes soft tissue, bone, and air. To suppress irrelevant structures and emphasize renal parenchyma and lesions, we clipped HU values to a fixed window (window width = 400, window level = 50), as is standard in abdominal imaging. This was followed by min–max normalization to rescale intensity values to the range [1,1]11[-1,1][ - 1 , 1 ], promoting training stability and improving convergence.
@@ 4.6 Metrics
4.6.1 Feature Extraction Evaluation
To evaluate the performance of the feature extraction model, we adopt standard metrics tailored to each task type. For classification tasks—such as detecting cysts, masses, or tumors—we report Accuracy, Precision, Recall, and F1-score. To account for class imbalance (e.g., the tumor class has very few positive samples), we additionally report the Area Under the Receiver Operating Characteristic Curve (AUC-ROC), which offers a more reliable measure of discriminatory performance in imbalanced scenarios. For regression tasks, such as predicting lesion size, we use Mean Squared Error (MSE) to quantify the average deviation between the predicted and ground truth sizes (in cm). This metric captures both systematic and random errors in continuous predictions.
@@ 4.6.2 Report Generation Evaluation
To assess the quality of the generated radiology report sentences, we report standard Natural Language Generation (NLG) metrics including BLEU-4 [33], ROUGE-1, ROUGE-2, and ROUGE-L [34]. These metrics capture n-gram overlap and longest common subsequence between the generated and reference sentences, providing a lexical-level assessment of fluency and content similarity.
@@ 5.1 Implementation Details
For the clinical feature prediction tasks, we performed grid search over multiple ResNet variants (ResNet-18, ResNet-50, and ResNet-101), initialized with ImageNet-pretrained weights. The input images were preprocessed using the augmentation strategies described in Section 4.5, including adjacent slice incorporation, intensity clipping, normalization, and resizing to $512\times 512512 × 512$ resolution. Training was conducted with a batch size of 16, a learning rate of 1e-4, and the Adam optimizer.
@@
We fine-tuned the Qwen2.5-VL-7B-Instruct model using LoRA (rank = 8) with the LLaMA-Factory framework. CT slices were preprocessed with fixed windowing (level = 50, width = 400) and resized to 512×512 pixels. Structured features were embedded into the prompt alongside the CT image (Appendix A). Training used a learning rate of 1e-4, batch size of 4 (with gradient accumulation = 8), and a cosine scheduler over 200 epochs. Mixed-precision training (bf16) was enabled, and evaluation was performed every 50 steps.
@@
All experiments were conducted on a single NVIDIA H100 B200 GPU with 192 GB of memory. Code was implemented in PyTorch 2.2 and Hugging Face Transformers, with all random seeds fixed for reproducibility.
5.2 Feature Extraction Results
@@
Table 2 summarizes the performance of our proposed feature extraction model compared to a random baseline across all annotated abnormality types.
The proposed model substantially outperforms the random baseline across all abnormality detection tasks, indicating reliable learning behavior and good generalization performance. For most binary features—such as Position, Cyst, and Tumor—the model achieved substantial improvements in AUC (e.g., 0.7063 vs. 0.3168 for Tumor) and F1-score (0.7400 vs. 0.3510), indicating accurate identification of clinically relevant findings. Particularly strong performance was observed for the Enhancement (F1 = 0.8859) and Tumor (F1 = 0.7400) categories, which are clinically important for lesion characterization and cancer diagnosis, respectively.
For the multi-class task of Attenuation, the model yielded moderate results, with a precision of 0.6300 and F1-score of 0.6167. This is likely due to overlapping Hounsfield Unit (HU) values among hypo-, iso-, and hyperattenuating tissues, as well as the limited availability of explicit attenuation descriptions in the source reports. Nevertheless, the model still achieved more than a twofold improvement in accuracy over the baseline (0.6827 vs. 0.3202), suggesting its ability to learn meaningful attenuation cues.
The evaluation results for the Exophytic label appear perfect, but are misleading. Since the dataset contains only a single instance labeled as “endophytic,” which appears in both the training and validation sets, the model likely memorizes this case. As a result, the reported performance does not reflect the true generalization ability for this feature.
For lesion size estimation, the model yielded an overall MSE of 0.9959 cm. Although this value may appear relatively high, further analysis revealed that the performance drop was primarily driven by a single cross-validation fold, in which the validation set contained several large lesions that were underrepresented in the corresponding training set. In other folds, the model demonstrated strong consistency and accuracy, suggesting that adequate size diversity in the training data can substantially enhance regression performance.
Table 2: Comparison of Model Performance Across Abnormality Features
@@
Proposed 0.9959
5.3 Report Generation Results
Table 3 presents the average performance of our report generation model across 5-fold cross-validation, evaluated using standard natural language generation (NLG) metrics. The model achieved a mean BLEU-4 score of 29.58, indicating moderate n-gram overlap with ground-truth sentences. ROUGE scores were also consistent, with ROUGE-1 at 28.31, ROUGE-2 at 9.59, and ROUGE-L at 21.99, reflecting reasonable lexical and structural alignment between the generated and reference sentences.
These results suggest that the model is capable of generating radiology sentences that are both clinically coherent and semantically relevant. However, the relatively low ROUGE-2 score indicates that bigram-level fluency and specificity remain limited—likely due to variability in radiology writing styles and the limited size of the training corpus. Incorporating domain-specific pretraining or reinforcement learning with clinical feedback may further improve factual accuracy and linguistic fidelity in future work.
Table 3: Performance of report generation
Metric BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
Mean Value 29.58 28.31 9.59 21.99
6 Limitations

### v2

@@ 3.2.3 CT Slice
2D slices were extracted from the 3D CT volumes based on explicit references in the radiology reports (e.g., “series 4 image 167 ”). Slice positions were determined by sorting the DICOM series using the SliceLocation attribute, followed by index-based matching to the referenced image number. Each slice was aligned with its corresponding sentence and feature labels to form complete data triplets: (report sentence, feature label, CT slice). We also standardized the spatial resolution of all CT slices by resizing each to a fixed dimension of $512\times 512$ pixels. For slices larger than this size, we applied center cropping; for smaller ones, we padded the surrounding region with zeros. This resizing strategy ensured consistency across inputs while preserving lesion scale and spatial integrity—particularly important for size-sensitive tasks. In addition to spatial normalization, we applied standard intensity preprocessing. Raw CT scans were stored in Hounsfield Units (HU), spanning a wide dynamic range that includes soft tissue, bone, and air. To suppress irrelevant structures and emphasize renal parenchyma and lesions, we clipped HU values to a fixed window (window width = 400, window level = 50), as is standard in abdominal imaging. All CT slices were manually reviewed to confirm alignment with the renal region. Slices that were incorrectly referenced or anatomically misaligned in the original report were excluded.
@@ 3.3 Data Splitting and Label Balance
A 5-fold cross-validation strategy was adopted to ensure robustness and generalizability. Data splitting was performed at the annotation level, where each sample corresponds to a unique triplet of (report sentence, label, CT slice). This approach enables stratification across a larger number of samples while preserving label diversity within each fold. To mitigate sampling bias, we applied stratified sampling [31, 32] based on renal feature distributions. For highly imbalanced features—such as the “Exophytic” category, which included only one “endophytic” case—we manually ensured that the minority class was represented in both training and validation sets. The detailed label distribution for a representative fold is provided in Appendix B.
@@ 3.4 Model
To improve clinical interpretability, we employ a two-stage modeling pipeline (Figure 1). In the first stage, clinical features are detected from 2D CT slices using dedicated computer vision models. In the second stage, the detected features—together with the corresponding CT slice—are input into a fine-tuned VLM to generate a report. This modular architecture supports both structured feature evaluation and flexible, context-aware report generation.
3.4.1 Feature Detection Model
We use a ResNet backbone [33] as the base encoder for CT slice-level clinical feature detection. Each renal feature label is treated as an independent task:
@@ 3.4.1 Feature Detection Model
The Exophytic and Tumor categories are highly imbalanced in our dataset (e.g., only one endophytic case among 130 lesions), making conventional supervised classification ineffective. This imbalance may partly arise from reporting conventions, as radiologists at our institution rarely use the term “endophytic” explicitly. For these features, we employ anomaly detection techniques [34] trained exclusively on negative examples.
@@ 3.4.2 Report Generation Model
We used Qwen-VL 2.5 [22] and LLaMA-3.2-Vision [35] as the backbone VLMs for radiology report generation. To assess the contribution of different input modalities, we evaluated three configurations: (1) structured clinical features only, (2) 2D CT slice images only, and (3) both modalities combined. Each model was tested in both fine-tuned (ft) and zero-shot (zs) settings, resulting in variants such as ft-image, ft-feature, and ft-both. Fine-tuning was conducted using LLAMA-Factory [36] with instruction-style supervision, where the input consists of a prompt (containing clinical features and/or an image) and the output is the corresponding expert-written report sentence. All experiments followed a consistent training and evaluation protocol. Specific prompt formats for each modality setting are detailed in Appendix A.
@@ 3.5 Data Augmentation
To improve model robustness and leverage spatial continuity in CT volumes, we applied augmentation strategies inspired by prior work [15]. Specifically, for each annotated 2D CT slice, we included its adjacent slices at offsets of 1-1 and +1+1, effectively tripling the training data and introducing slight anatomical variation without altering underlying pathology. For the feature detection model, we also experimented with different window level settings to enhance contrast and highlight renal abnormalities. After intensity clipping, voxel values were rescaled to the range [1,1][-1,1] using min-max normalization to promote training stability and improve convergence.
@@ 3.6 Metrics
3.6.1 Clinical Feature Evaluation
To evaluate the performance of the feature extraction model, we adopt standard metrics tailored to each task type. For classification tasks—such as detecting cysts, masses, or tumors—we report Accuracy, Precision, Recall, and F1-score. To account for class imbalance (e.g., the tumor class has very few positive samples), we additionally report the Area Under the Receiver Operating Characteristic Curve (AUC-ROC), which offers a more reliable measure of discriminatory performance in imbalanced scenarios. For regression tasks, such as predicting lesion size, we use Mean Squared Error (MSE) to quantify the average deviation between predicted and ground truth sizes (in cm). This metric captures both systematic and random errors in continuous predictions.
@@ 3.6.2 Report Generation Evaluation
To evaluate the similarity of generated radiology report sentences, we used three standard NLG metrics: BLEU [37], ROUGE-L [38], and METEOR [39]. These metrics assess lexical similarity between generated and reference sentences by capturing n-gram overlap, longest common subsequence, and semantic alignment with synonym matching, respectively. In practice, we observe that different implementations of these metrics—such as those provided by the COCO Caption toolkit [40], NLTK [41], and Hugging Face—can yield slightly different numerical values due to variations in preprocessing, tokenization, and smoothing strategies. To ensure consistency and reproducibility, we adopt the implementations provided by the Hugging Face evaluate library [42] in all experiments.
@@ 4.1 Implementation Details
For clinical feature detection, we performed a grid search across multiple ResNet variants (ResNet-18, ResNet-50, and ResNet-101), each initialized with ImageNet-pretrained weights. Input images were preprocessed using the augmentation strategies described in Section 3.5, including adjacent slice incorporation, intensity windowing, normalization, and resizing to a fixed resolution of $512\times 512$ pixels. Models were trained with a batch size of 16, a learning rate of 1e-4, and the Adam optimizer. The best checkpoint was selected based on the highest validation F1 score.
@@
For report generation, we fine-tuned the Qwen2.5-VL and LLaMA3.2-Vision models using Low-Rank Adaptation (LoRA) with rank = 8. Structured clinical features were embedded into the input prompt alongside the CT slice image (see Appendix A). Training was performed with a learning rate of 1e-4, batch size of 4 (gradient accumulation = 8), and a cosine learning rate scheduler over 20 epochs. Mixed-precision training (bf16) was enabled, and evaluation was conducted every 20 steps. The optimal checkpoint was selected based on the highest ROUGE-L score on the validation set.
@@
All experiments were conducted on a single NVIDIA H100 B200 GPU with 192 GB of memory. The implementation was based on PyTorch 2.2 and Hugging Face Transformers, with all random seeds fixed for reproducibility.
4.2 Clinical Feature Detection Results
Table 2 summarizes the performance of our clinical feature detection model compared to a random baseline across all annotated abnormality types. The proposed model consistently outperformed the random baseline across all detection tasks. For most binary features,the model achieved gains in both AUC (e.g., 0.7063 vs. 0.3168 for Tumor) and F1 score (0.7400 vs. 0.3510). Performance was particularly high for Enhancement (F1 = 0.8859) and Tumor (F1 = 0.7400). For the multi-class classification task of Attenuation, the model achieved moderate results, with a precision of 0.6300 and an F1 score of 0.6167. Although the evaluation metrics for the Exophytic label appear perfect, they are misleading. The dataset contains only one instance labeled as “endophytic”, which appears in both the training and validation sets. For lesion size estimation, the model reported a mean squared error (MSE) of 0.9959 cm. Further analysis revealed that the error was primarily driven by a single cross-validation fold containing several unusually large lesions that were underrepresented in the training set. In other folds, the model exhibited strong consistency and accuracy.
@@ 4.2 Clinical Feature Detection Results
Table 2: Performance comparison between the proposed clinical feature detection model and a random baseline across all annotated features.
@@
Proposed 0.9595
4.3 Radiology Report Generation Results
We evaluated the quality of generated radiology reports using both NLG metrics and clinical feature accuracy. NLG metrics compare generated reports against radiologist-written reference sentences by measuring lexical similarity (Table 3). To evaluate clinical correctness, we automatically extracted structured features from generated reports using an LLM-based parser and compared them with ground truth labels. Because the generated reports were produced by an LLM, they exhibit consistent formatting, reduced linguistic variability, and structured presentation of clinical findings, which facilitated information extraction with an accuracy exceeding 98%. This evaluation captures whether the generated text faithfully conveys key clinical findings (Table 4). All results are reported as averages over five cross-validation.
4.3.1 Textual Evaluation of Generated Report
For fine-tuned Qwen models, NLG scores remained relatively consistent across different input settings (clinical features only, image only, or both). In the zero-shot setting, models tend to produce full-length radiology reports (see Figure 3), which differ in structure and verbosity from concise, expert-authored references, resulting in lower NLG scores. We also observed that smaller models, such as LLaMA-11B, failed to converge under our experimental settings. Effective adaptation required scaling to the 90B-parameter variant.
4.3.2 Generated Clinical Feature Evaluation
Since clinical features were extracted directly from the generated reports, no confidence probabilities were available, making AUC and other thresold-based metrics inapplicable. From Table 4, we observed that models relying solely on image input (e.g., ft-image, zs-image variants) showed reduced performance for several features, including Position, Exophytic, Attenuation, and Enhancement. Interestingly, fine-tuned models often yielded lower clinical feature accuracy than their zero-shot counterparts. Fine-tuned models exhibited more frequent hallucinations for features such as Cyst, Mass, and Tumor, where as zero-shot models were comparatively less affected. Regarding the Size_cm feature, evaluated using MSE, we observe no consistent correlation between model scale and performance. Even larger models like Qwen-32B or LLaMA3-90B failed to outperform smaller ones. In some cases, models generated implausible multi-dimensional size estimates despite receiving only 2D inputs.
4.4 Case Study
To better understand model behavior, we examined two representative cases (Figure 3). Overall, fine-tuned models tend to generate reports that are structurally closer to those written by radiologists. In contrast, the zero-shot model (Qwen-32Bzs-both) produced lengthy, well-structured reports that resemble radiology templates. This observation aligns with the trends in NLG evaluation metrics. Interestingly, the zero-shot model closely adhered to the input prompt, reproducing the provided information faithfully without extrapolating beyond it. On the other hand, fine-tuned models occasionally introduced new content not explicitly mentioned in the prompt. Some of this added content is correct, while some are factually inaccurate. For example, in Case 2, Qwen-32Bft-both correctly inferred that the lesion is located in the interpolar region, which matches the ground truth. However, it also hallucinates a three-dimensional lesion size, despite only the largest diameter being provided in the prompt. Moreover, we observed that the overall quality and factual correctness of generated reports heavily depended on the accuracy of the input prompt. When models relied solely on the image (e.g., Qwen-32Bft-image), the generated reports often deviated more significantly from the ground truth.
Table 3: Comparison of lexical performance in radiology report generation across model configurations and input modalities. “Qwen” refers to the Qwen2.5-VL series; “LLaMA” refers to the LLaMA-3.2-Vision series.
Model Clinical Feature 2D Slice Image BLEU-1 BLEU-4 ROUGE-L METEOR
Qwen-32Bft-both 0.299 0.079 0.281 0.325
Qwen-32Bft-image 0.293 0.082 0.253 0.323
Qwen-32Bft-feature 0.278 0.069 0.273 0.309
Qwen-32Bzs-both 0.020 0.003 0.091 0.060
Qwen-32Bzs-feature 0.018 0.003 0.043 0.089
Qwen-32Bzs-image 0.014 0.002 0.033 0.075
Qwen-7Bft-both 0.271 0.069 0.275 0.307
Qwen-7Bft-feature 0.294 0.087 0.280 0.309
Qwen-7Bft-image 0.306 0.079 0.275 0.330
Qwen-7Bzs-both 0.031 0.004 0.128 0.078
LLaMA-11Bft-both 0.001 0.000 0.017 0.101
LLaMA-90Bft-both 0.269 0.054 0.253 0.286
Note: Model names use the prefix “ft-” for fine-tuned and “zs-” for zero-shot. The suffixes “image”, “feature”, and “both” indicate which modality/modalities are included in the input prompt.
Table 4: Evaluation of clinical feature performance extracted from generated reports.
Model Feature Accuracy Precision Recall F1 / MSE
Qwen-32Bft-both Position 0.7175 0.7330 0.7262 0.7132
Exophytic 0.6133 0.6000 0.4471 0.5033
Attenuation 0.4157 0.1600 0.2735 0.1998
Enhancement 0.5772 0.5583 0.3864 0.4503
Cyst 0.5687 0.5982 0.5974 0.5608
Mass 0.4541 0.5463 0.5690 0.3939
Tumor 0.9392 0.5767 0.5919 0.5841
Size_cm 1.7444
Qwen-32Bft-feature Position 0.6938 0.7268 0.7041 0.6973
Exophytic 0.7367 0.7000 0.6007 0.6419
Attenuation 0.4794 0.2792 0.3098 0.2543
Enhancement 0.5594 0.5033 0.4114 0.4441
Cyst 0.4452 0.5371 0.5144 0.4229
Mass 0.3207 0.5324 0.5269 0.3066
Tumor 0.8333 0.4696 0.4400 0.4533
Size_cm 1.1081
Qwen-32Bft-image Position 0.4487 0.4494 0.4420 0.4267
Exophytic 0.1833 0.4000 0.1086 0.1683
Attenuation 0.2946 0.3176 0.2061 0.2448
Enhancement 0.1244 0.4000 0.1043 0.1583
Cyst 0.4909 0.5377 0.5327 0.4853
Mass 0.3735 0.5094 0.4981 0.3436
Tumor 0.9310 0.4728 0.4917 0.4820
Size_cm 1.9741
Qwen-32Bzs-both Position 0.7175 0.7330 0.7262 0.7132
Exophytic 1.0000 1.0000 1.0000 1.0000
Attenuation 0.5136 0.1712 0.3333 0.2254
Enhancement 0.7989 0.6300 0.6281 0.6154
Cyst 0.7658 0.7771 0.7482 0.7403
Mass 0.8702 0.6548 0.6113 0.6038
Tumor 0.9620 0.7344 0.7460 0.7400
Size_cm 1.0987
Qwen-32Bzs-feature Position 0.7175 0.7330 0.7262 0.7132
Exophytic 1.0000 1.0000 1.0000 1.0000
Attenuation 0.5136 0.1712 0.3333 0.2254
Enhancement 0.8389 0.6449 0.6074 0.5993
Cyst 0.7735 0.7848 0.7573 0.7494
Mass 0.8862 0.7448 0.6350 0.6592
Tumor 0.9469 0.6137 0.6420 0.6259
Size_cm 1.0987
Table 5: Evaluation of clinical feature performance extracted from generated reports. (continued)
Model Feature Accuracy Precision Recall F1 / MSE
Qwen-32Bzs-image Position 0.5040 0.5207 0.4990 0.4962
Exophytic 0.1667 0.2667 0.1000 0.1405
Attenuation 0.4289 0.3090 0.2862 0.2919
Enhancement 0.2483 0.3275 0.2667 0.2486
Cyst 0.5147 0.4803 0.4918 0.4739
Mass 0.5097 0.4654 0.4083 0.3883
Tumor 0.8925 0.4719 0.4717 0.4709
Size_cm 3.1239
Qwen-7Bft-both Position 0.7175 0.7330 0.7262 0.7132
Exophytic 0.6533 0.7800 0.6271 0.6833
Attenuation 0.4088 0.3041 0.2696 0.2361
Enhancement 0.5844 0.5875 0.4083 0.4596
Cyst 0.5445 0.6361 0.6064 0.5218
Mass 0.3340 0.5563 0.5772 0.3190
Tumor 0.9307 0.4726 0.4913 0.4817
Size_cm 1.2672
Qwen-7Bft-feature Position 0.6921 0.7346 0.6980 0.7020
Exophytic 0.8683 0.7000 0.6857 0.6923
Attenuation 0.4580 0.3274 0.3042 0.2739
Enhancement 0.6467 0.5124 0.4357 0.4667
Cyst 0.5648 0.6045 0.6054 0.5575
Mass 0.4509 0.5704 0.6732 0.4142
Tumor 0.8929 0.4716 0.4720 0.4715
Size_cm 1.8435
Qwen-7Bft-image Position 0.4504 0.4433 0.4403 0.4355
Exophytic 0.4467 0.5333 0.4329 0.4421
Attenuation 0.3488 0.3244 0.2482 0.2679
Enhancement 0.1022 0.3000 0.0650 0.1067
Cyst 0.4961 0.5294 0.5295 0.4937
Mass 0.3942 0.4845 0.4609 0.3404
Tumor 0.9170 0.4727 0.4846 0.4782
Size_cm 1.9127
Qwen-7Bzs-both Position 0.7175 0.7330 0.7262 0.7132
Exophytic 1.0000 1.0000 1.0000 1.0000
Attenuation 0.5136 0.1712 0.3333 0.2254
Enhancement 0.9083 0.7258 0.6717 0.6859
Cyst 0.7496 0.7656 0.7139 0.7182
Mass 0.8787 0.7090 0.6886 0.6846
Tumor 0.9395 0.6591 0.7340 0.6836
Size_cm 1.0987
Table 6: Evaluation of clinical feature performance extracted from generated reports. (continued)
Model Feature Accuracy Precision Recall F1 / MSE
LLaMA3-11Bft-both
Position 0.5508 0.6400 0.5325 0.5549
Exophytic 0.4517 0.3417 0.2743 0.2941
Attenuation 0.1982 0.1866 0.1567 0.1398
Enhancement 0.3356 0.4014 0.2414 0.2681
Cyst 0.4247 0.5090 0.5002 0.3632
Mass 0.2468 0.4442 0.4098 0.2307
Tumor 0.9458 0.4729 0.5000 0.4860
Size_cm 4.3634
LLaMA3-90Bft-both Position 0.7175 0.7330 0.7262 0.7132
Exophytic 0.9333 0.8000 0.8000 0.8000
Attenuation 0.4686 0.1679 0.3019 0.2151
Enhancement 0.7783 0.6258 0.5650 0.5809
Cyst 0.5688 0.6092 0.6048 0.5412
Mass 0.3815 0.5613 0.6027 0.3501
Tumor 0.9229 0.5758 0.5830 0.5792
Size_cm 1.0864
Figure 3: Comparison of model-generated radiology reports and extracted features under different input modalities (case 1).
Figure 4: Comparison of model-generated radiology reports and corresponding extracted features across different input modalities (case 2).
5 Discussion
5.1 Clinical Feature Detection
Overall, the proposed model consistently outperformed the random baseline across all feature detection tasks, demonstrating robust learning and strong generalization. For most binary features—such as Position, Cyst, and Tumor—the model achieved notable gains in both AUC and F1 score, showing its effectiveness in identifying clinically relevant findings. Performance was particularly strong for the Enhancement and Tumor categories, both of which are critical for lesion characterization and cancer diagnosis.
By contrast, the moderate performance on Attenuation highlights the difficulty of this task. This limitation is likely due to overlapping HU values among hypo-, iso-, and hyperattenuating tissues, combined with the scarcity of explicit attenuation descriptions in the source reports. Nevertheless, the model achieved more than a twofold improvement in accuracy compared to the baseline, suggesting that it effectively captures meaningful attenuation patterns despite data constraints. The apparently perfect metrics for the Exophytic category are misleading. With only one “endophytic” instance present across training and validation, this likely caused overfitting via memorization, meaning the reported metrics do not reflect true generalization for this feature. For lesion size estimation, further analysis revealed that one fold showed substantially higher error than the others. This was primarily due to the presence of several unusually large lesions in the validation set, which were underrepresented in the corresponding training data. A more balanced distribution of lesion sizes in the training set could therefore improve regression performance.
5.2 Radiology Report Generation
The results highlight several important trends in report generation. First, models relying solely on image input struggled to capture detailed clinical findings. When models relied only on the image (e.g., Qwen-32Bft-image), the generated reports deviated more significantly from the ground truth. This suggests that our two-stage framework—first extracting structured clinical features and then generating the report—is more reliable than directly generating text from images, particularly in terms of clinical fidelity. Second, fine-tuned models often underperformed their zero-shot counterparts in clinical feature accuracy. We hypothesize that this may be due to fine-tuning on ground-truth reports, which encourages the model to imitate the linguistic style of radiologists. Such stylistic imitation can lead to the omission of explicit clinical details or the reproduction of incorrect information, thereby negatively affecting the expression of key clinical features. Moreover, the hallucinating in fine-tuned models may arise from memorized report templates that promote overgeneralization and misrepresentation of clinical content. Zero-shot models were comparatively less affected by this problem. Lesion size estimation remains a particularly challenging task. In several cases, models generated 3-dimensional size despite receiving only inputs 2-dimensional size. Finally, no consistent relationship was observed between model scale and clinical accuracy. This may reflect the limited size of our dataset, which restricts the benefits of scaling to larger models. Overall, these findings suggest that while large vision–language models are capable of producing fluent reports, challenges remain in ensuring factual correctness and clinical fidelity.
5.3 Limitations